Journalist Report – 12 April 2018

Author: Victoria Varone

Today was one of those days that highlights just how passionate, adaptable, and hard-working space dreamers really are.

In the face of weather that disrupted our EVA plans, the whole crew stepped up to the plate, offering up an unbelievable variety of skills and ideas to make repairs and adjustments with a limited supply of tools and hardware. They pooled their backgrounds, areas of expertise, physical capabilities, and improvisation skills to come up with a solution to the problem at hand.

Nothing brings out the strength of such dreamers more than a problem that needs to be solved, be it an Earthly issue or a Mars mission. I’m incredibly proud of this crew today and never except anything less than amazing when we’re together.

In honor of Yuri’s Night: “When I orbited the Earth in a spaceship, I saw for the first time how beautiful our planet is. Mankind, let us preserve and increase this beauty, and not destroy it!”

GreenHab Report – 12 April 2018

Name: Andreea Radulescu
Date: 12 Apr 2018

Environmental control:

Ambient with door open
Shade cloth on
Fan not on.

Average temperatures: 39.7℃
Low 14.2℃
High 53.2℃

Hours of supplemental light: None

Changes to crops: Was able to pick a lot of tomatoes.

Daily water usage for crops: 5 gal

Time(s) of watering for crops: Once in the morning around 10:00 am

Moringa research observations: None

Changes to research plants: No

Aquaponics: None

Narrative: Picked few more tomatoes for the team which we will consume at dinner.

Support/supplies needed: None

EVA Report – 12 April 2018

EVA #6 – 12Apr2018 – Sol 4

Crew members: Andreaa Radulescu, Ashok Narayanamoorthi, Richard Blakeman

Objectives:
Scout for good sites to perform medical emergency simulations.

Operations:
0845 Start EVA
0900 Surveyed URC North Site
0930 Returned to HAB at request of commander
1000 EVA end, transfer to engineering EVA.

Narrative:
This EVA was meant for scouting sites for a medical emergency situation. One site was surveyed before the crew members on the EVA were recalled to help with a repair in the HAB from damage caused by severe winds.

Vehicles
1. Blue #1
2. Blue #2
3. Blue #3
_______________

EVA #7 – 12Apr2018 – Sol 4

Crew members: Richard Blakeman, Ashok Narayanamoorthi, Joe Dituri, Andreaa Radulescu, Victoria Varone

Objectives:
– All crew transfer to the RAM and clean and sanitize the water tank.
– Garbage pails tied to trailer as they were blowing away in the strong wind.

Operations:
1630 – 1745

Narrative:
Today’s second EVA was a productive engineering EVA. The tank is 85% cleaned and we intend to complete further final cleaning tomorrow.

Vehicles
– None

Operations Report – 12 April 2018

SOL 4

Name of person filing report: Joe Dituri

Non-nominal systems: 1. The main hab door broke during the wind.
2. The circuit breaker tripped.

Notes on non-nominal systems: During severe winds again today, the door swung open HARD and almost came off its hinge. We’ve performed a sturdy and strong repair that is holding beautifully so far.

After it became very cold in the hab, we checked the circuit breaker and noticed it had tripped. It was reset.

Generator (hours run): 13 hours

Solar ‒ SOC 98%

Diesel – 75% fuel

Propane – 66% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 960 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133383.5

Toilet tank emptied: Yes. Last emptied April 11, 2018.

ATV’s Used: Blue ATVs #1, #2, and #3

Oil Added? No But checked and OK.

ATV Fuel Used: 10%

# Hours the ATVs were Used today: 2

Notes on ATVs: All working well and ready for use.

Deimos rover used: No
Hours: 119.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Spirit rover used: No
Hours: 31
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Crew is doing well, we had to stay on alert today for the strong winds. Will remain alert for continued winds tomorrow.

Summary of internet: 198MB at 1930.

Summary of suits and radios: Suits all charging, 4 one-piece suits and 1 two piece suit used today, radios working fine.

Summary of Hab operations: Repaired main hab door latch and wooden rail after strong winds blew the door all the way out, broke the latch, and knocked the rail off the stairs. Cleaned and decontaminated the trailer water tank in the RAM. The tank requires further cleaning.

Summary of GreenHab operations: See GreenHab report.

Summary of ScienceDome operations: None

Summary of RAM operations: The trailer water tank was cleaned in the RAM.

Sol Summary Report – 12 April 2018

Sol 4

Summary Title: Whatever It Takes

Author: Victoria Varone

Mission Status:

Sol Activity Summary: It was quite a busy day for Crew 192 as we spent almost all of our time literally “battening down the hatches” as severe winds caused the HAB some physical wear that needed repair and periodic observation.

It began at ~0220 when the wind peeked. We had to tie down the center wooden cap for the structure because it kept slamming against the ceiling and being pulled up by the wind. It is now secure.

We had an approved short morning scouting EVA …The commander recalled the team when the front airlock door vibrated and broke free from its hasp and the door swung open so hard that it broke the banister and damaged the habitat. The commander decided to have the team emergent repair the HAB by reattaching the hasp and screwing the banister back in place. The external repairs to the HAB must be made at a later date because they were not emergent and beyond the ability of the crew.

We completed another engineering EVA (ALL HANDS EFFORT) to clean the trailer water tank in the RAM so that our next crew can maintain a large and safe supply of water. This involved stringing 550 cord to make a safety line which allowed all personnel to get to the HAB without issue in the mounting wind.

Look Ahead Plan: The strong winds are forecast to continue into tomorrow and should slow down after that. Once it does, we will resume our exploration EVAs, continually exploring new sites and familiarizing ourselves further with ones we’ve already seen.

Anomalies in Work: None

Weather: Chilly and extremely windy. Winds blowing at approximately 30 mph with gusts 46 mph.

Crew Physical Status: All crew are physically healthy and safe.